OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = x - 8. y = x + 7 y = x + 5 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

y=mx+c, m=1 now choose a point (x,y) and imagine it is on the same line with (-2,6). NOW JUST find the gradient using your points and equate it to 1

OpenStudy (compassionate):

\(\LARGE\color{black}{{Hello!~:)}}\) Have you found your answer?

OpenStudy (anonymous):

Nope....

OpenStudy (anonymous):

Im lost with this whole algebra stuff since I dont get the difference between the equations since i dont understand it

OpenStudy (compassionate):

Okay, well, hello, my name is Stephen. You can call me Steve if you'd like. "Choose the equation of the line passing through the point (-2, 6) and parallel to y = x - 8." Okay, we know the line is parallel. Remember that parallel lines have the same slope. If the equation is in y = mx + b form, were m is the slope. Then: y = x - 8, your slope is: x, or just 1. We are also given the points (-2, 6) Info so far Slope: 1 Points: (-2, 6) We can use the point-slope form to find the equation. y - y1 = m(x - x1) Plug in the terms y - 6 = 1(x - [2)] Then we solve the equation y - 6 = x - 2 y = x - 2 + 6 Can you solve it from there?

OpenStudy (anonymous):

Wait so x is x and y is y and m - slope right?

OpenStudy (anonymous):

Yes, and how you can remember if a problem is an equation is it always has an equals sign in the middle of the numbers.

OpenStudy (anonymous):

Ok thanks!! I never really understood that. .